Nov 15, 14:32 UTC
Services have been fully operational for the last 3 hours. This incident is resolved.
Nov 15, 13:09 UTC
We have fully recovered the availability of all services. All affected environments should now be working as expected. We're monitoring the situation closely.
Nov 15, 09:47 UTC
Data reallocation is still ongoing, affecting some environments. Affected environments can experience reduce input/output performance.
Nov 15, 07:57 UTC
We are continuing to monitor the data reallocation process. Changes have been implemented to reduce the performance impact of the ongoing procedures, but the region will continue experiencing degraded performance until this is complete. We currently estimate that the ongoing process may still take a number of hours to complete fully.
Nov 15, 05:18 UTC
Routine Ceph OSDs being added into the region is currently resulting in an elevated level of Input/Output (IO) and now throttling above normal levels.
Nov 15, 04:24 UTC
We have detected an issue affecting service on the de-2.platform.sh region. Our Operations team has been notified and is currently working to restore service. Projects on affected regions may experience degraded performance.
We will update you as soon as we have further information.
Nov 15, 04:17 UTC